The mass of a spaceship is 1000 kg. It is to be launched from the earth's surface out into free space. The value of 'g' and 'R' (radius of earth) are 10 m/s2 and 6400km respectively. The required energy for this work will be ; (1) 6.4 x 1011 Joules (2) 6.4 x 108 Joules (3) 6.4 x 109 Joules (4) 6.4 x 1010 Joules ?

Grade:11

2 Answers

BHAVISHYA UPADHYAY
6 years ago
answer is 64×10^10jouleW=0-GMm/R=GMm/R=gR^2×m/R=1000×10×6400×10^3=64×10^9if you have further any queries here ask
Gowtham
6 years ago
=GMm/rThe required energy for this work will be
we know that force = GMm/R^2
mg=GMm/R^2
g=GM/R^2
mgR=GMm/R